# Content-based image retrieval for multi-class volumetric radiology images: a benchmark study

### v1

@@ 4.3 Embeddings
It was shown that embeddings generated from self-supervised models are slightly better for image retrieval tasks than those derived from regular supervised models. This is true for coarse anatomical regions with 29 labels (see LABEL:tab:recall-29-regions) as well as fine-granular anatomical regions with 104 regions (see LABEL:tab:recall-104-regions). This is roughly preserved for all modes of retrieval (i.e. slice-wise, volume-based, and region-based). More generally, the differences in recall across differently pre-trained models (except pre-trained from fractal image) are very small and except for slice-wise close to $1.0$. Practically, the exact choice of the feature extractor should not be noticeable to a potential user in a downstream application. Further, it can be concluded that pre-training on general natural images (i.e. ImageNet) resulted in slightly more performant embedding vectors than domain-specific natural images (i.e. RadImageNet). This is somewhat surprising and is subject to further research.
Although, the model pre-trained of formula-derived synthetic images of fractals (i.e. Fractaldb) showed the lowest recall accuracy the absolute values are surprisingly high considering that the model learned visual primitives out of rendered fractals. This is very encouraging as the Formular-Driven Supervised Learning (FDSL) can easily be extended to the very high number of data points per class and also several virtual classes within one family of formulas (Kataoka et al., 2022). Additionally, the mathematical space of formulas for producing visual primitives is virtually infinite and thus it is the subject of further research whether radiology-specific visual primitives can be created that outperform natural image-based pre-training. Again, FDSL does not require the effort of data collection, curation, and annotation. It can scale to a large number of samples and classes which potentially results in a very smooth and evenly covered latent space.
Embeddings derived from DreamSim architecture showed the highest overall retrieval recall in the region-based retrieval with re-ranking. This was true for 29 coarse anatomical regions as well as the original 104 fine-granular regions as used in TotalSegmentator. DreamSim is an ensemble architecture that uses multiple ViT embeddings with additional fine-tuning using synthetic images. It is plausible that an ensemble approach outperforms single-architecture embeddings (i.e. DINOv1, DINOv2, SwinTransformer, and ResNet50). Therefore, the usage of DreamSim is the currently preferred method of embedding generation.
@@
slice-wise .$839\pm.156$$819\pm.148$$\mathbf{.849\pm.148}$$824\pm.146$$740\pm.167$$802\pm.14$
volume-based .$949\pm.072$$932\pm.064$$936\pm.063$$932\pm.067$$939\pm.078$$\mathbf{.952\pm.04}$
@@
region-based .$977\pm.047$$976\pm.051$$\mathbf{.979\pm.037}$$973\pm.042$$976\pm.033$$978\pm.033$
region-based re-ranked .$979\pm.045$$977\pm.050$$\mathbf{.987\pm.027}$$977\pm.041$$981\pm.031$$979\pm.04$
@@
Figure 9: An overview of three retrieval methods: (a) Slice-wise, where the retrieval is based on one selected slice e.g., the user zooms to a slice and retrieves the most similar slice (b) volume-based, where the retrieval is based on a complete volume, e.g., the user would like to retrieve similar volumes to the volumes under examination or simply filter the database (c) region-based, where the retrieval is based on the selected organ (or sub-volume), e.g, the user zooms in to a specific region and the most similar volume containing that region is retrieved.
4.4 Volume-based vs Region-based Retrieval
Since multiple organs (i.e. labels) are present in each query volume, there are essentially two ways in which image retrieval can be performed. The preferred choice depends on the context of the retrieval task in the real world. If the goal is to find a scan out of a database that is most similar to a complete query scan with the entirety of all present organs (think scan-id to scan-id but visual), then volume-based retrieval is the right choice. In contrast, if the experimenter is interested in a particular organ and its most similar counterpart in the database (and all other organs just happen to be in the same scan due to proximity), then region-based retrieval is advised. The second query scenario offers a higher retrieval recall. The slice-wise retrieval can find the most similar slice of a volume regardless of other slices. This is not usually a practical choice in real scenarios. Figure 9 visualizes the options.
@@ 5 Conclusion
Our study establishes a new benchmark for the retrieval of anatomical structures within 3D medical volumes, utilizing the TotalSegmentator dataset to facilitate targeted queries of volumes or sub-volumes for specific anatomical structures. The results highlight the potential of leveraging pre-trained embeddings, originally trained on non-medical images, for medical image retrieval across diverse anatomical regions with a wide size range.

### v2

@@ 4.3 Embeddings
It was shown that embeddings generated from self-supervised models are slightly better for image retrieval tasks than those derived from regular supervised models. This is true for coarse anatomical regions with 29 labels (see LABEL:tab:recall-29-regions) as well as fine-granular anatomical regions with 104 regions (see LABEL:tab:recall-104-regions). This is roughly preserved for all modes of retrieval (i.e. slice-wise, volume-based, region-based, and localized retrieval). More generally, the differences in recall across differently pre-trained models (except pre-trained from fractal image) are very small.Practically, the exact choice of the feature extractor should not be noticeable to a potential user in a downstream application. Further, it can be concluded that pre-training on general natural images (i.e. ImageNet) resulted in slightly more performant embedding vectors than domain-specific images (i.e. RadImageNet). This is unexpected and subject to further research.
Although, the model pre-trained of formula-derived synthetic images of fractals (i.e. Fractaldb) showed the lowest recall accuracy the absolute values are surprisingly high considering that the model learned visual primitives out of rendered fractals. This is very encouraging as the Formular-Driven Supervised Learning (FDSL) can easily be extended to very high number of data points per class and also several virtual classes within one family of formulas (Kataoka et al., 2022). Additionally, the mathematical space of formulas for producing visual primitives is virtually infinite and thus it is the subject of further research whether radiology-specific visual primitives can be created that outperform natural image-based pre-training. Again, FDSL does not require the effort of data collection, curation, and annotation. It can scale to a large number of samples and classes which potentially results in a very smooth and evenly covered latent space.
Embeddings derived from DreamSim architecture showed the highest overall retrieval recall in region-based and localized evaluations. DreamSim is an ensemble architecture that uses multiple ViT embeddings with additional fine-tuning using synthetic images. It is plausible that an ensemble approach outperforms single-architecture embeddings (i.e. DINOv1, DINOv2, SwinTransformer, and ResNet50). Therefore, the usage of DreamSim is currently the preferred method of embedding generation.
@@
slice-wise .$855\pm.108$$832\pm.118$$\mathbf{.863\pm.107}$$837\pm.114$$51\pm.52$$813\pm.124$
volume-based .$949\pm.072$$932\pm.064$$936\pm.063$$932\pm.067$$939\pm.078$$\mathbf{.952\pm.043}$
@@
region-based .$977\pm.047$$976\pm.051$$\mathbf{.979\pm.037}$$973\pm.042$$976\pm.033$$978\pm.039$
region-based re-ranked .$979\pm.045$$977\pm.050$$\mathbf{.987\pm.027}$$977\pm.041$$981\pm.031$$979\pm.045$
localized .$932\pm.105$$\mathbf{.941\pm.077}$$936\pm.089$$922\pm.099$$884\pm.111$$908\pm.099$
localized re-ranked .$944\pm.099$$955\pm.075$$\mathbf{.955\pm.062}$$943\pm.079$$907\pm.099$$947\pm.076$
@@
localized .$\mathbf{.929\pm.085}$$917\pm.091$$924\pm.078$$906\pm.092$$840\pm.145$$866\pm.118$
localized re-ranked .$944\pm.075$$951\pm.071$$\mathbf{.956\pm.055}$$940\pm.066$$890\pm.105$$935\pm.070$
Figure 11: An overview of three retrieval methods: (a) Slice-wise, where the retrieval is based on one selected slice e.g., the user zooms to a slice and retrieves the most similar slice (b) volume-based, where the retrieval is based on a complete volume, e.g., the user would like to retrieve similar volumes to the volumes under examination or simply filter the database (c) region-based, where the retrieval is based on the selected organ (or sub-volume), e.g, the user zooms in to a specific region and the most similar volume containing that region is retrieved, (d) localized, where the retrieval is based on the selected organ (or sub-volume) but the region in the retrieved volume is localized to the desired organ (or sub-volume), e.g, the user zooms in to a specific region and the system returns the localized region in the retrieved volume.
4.4 Volume-based, Region-based and Localized Retrieval
Since multiple organs (i.e. labels) are present in each query volume, there are essentially different ways in which image retrieval can be performed. The preferred choice depends on the context of the retrieval task in the real world. If the goal is to find a scan out of a database that is most similar to a complete query scan with the entirety of all present organs (think scan-id to scan-id but visual), then volume-based retrieval is the right choice. In contrast, if the experimenter is interested in a particular organ and its most similar counterpart in the database (and all other organs just happen to be in the same scan due to proximity), then region-based retrieval or localized retrieval is advised. Slice-wise retrieval can find the most similar slice of a volume regardless of the number of other slices. This is not usually a practical choice in real scenarios. Figure 11 visualizes the options.
4.5 Localization-ratio
LABEL:tab:localization-ratio-summary shows the average localization-ratio for with and without re-ranking ($L=15$). There is a drop for both 29 coarse and 104 original TS classes after re-ranking. However, both with and without re-ranking the localization-ratio is high, indicating that most of the slices contributing to the final retrieval of the volumes actually contain the desired anatomical region. However, based on (7) the choice of $L$ can impact this measure. Figure 12 shows localization-ratio for different values of $L$ as shown increasing the $L$ to more that $L=15$ decreases the localization-ratio. This indicates that the first highest values in vector $m_{SIM}$ based on (6) points to the exact slices that contain the desired anatomical structure. To improve the re-ranking localization-ratio, further studies can focus on optimizing the $L$ value based on heuristic, organ size, etc.
Figure 12: Average localization-ratio for different L values for re-ranking evaluation using DreamSim model (best-performing embedding) based on (7).
Table 22: Summary of the average localization-ratio and standard deviation between classes for 29 and 104 anatomical regions for $L=15$, the boldfaced values highlight the highest localization-ratio across feature extractors.
Model DINOv1 DINOv2 DreamSim SwinTrans. ResNet50
localization-ratio (29 anatomical regions) $0.842\pm 0.161$ $0.834\pm 0.144$ $\mathbf{0.864\pm 0.145}$ $0.835\pm 0.155$ $0.800\pm 0.168$ $0.815\pm 0.152$
localization-ratio (104 anatomical regions) $0.793\pm 0.140$ $0.763\pm 0.148$ $\mathbf{0.803\pm 0.130}$ $0.773\pm 0.143$ $0.722\pm 0.178$ $0.736\pm 0.162$
localization-ratio re-ranked (29 anatomical regions) $0.823\pm 0.169$ $0.803\pm 0.157$ $\mathbf{0.837\pm 0.159}$ $0.792\pm 0.164$ $0.727\pm 0.195$ $0.787\pm 0.171$
localization-ratio re-ranked (104 anatomical regions) $0.764\pm 0.148$ $0.754\pm 0.150$ $\mathbf{0.790\pm 0.142}$ $0.736\pm 0.145$ $0.662\pm 0.171$ $0.720\pm 0.166$
@@ 5 Conclusion
Our study establishes a new benchmark for the retrieval of anatomical structures within 3D medical volumes, utilizing the TotalSegmentator dataset to facilitate targeted queries of volumes or sub-volumes for specific anatomical structures. The results highlight the potential of leveraging pre-trained vision embeddings, originally trained on non-medical images, for medical image retrieval across diverse anatomical regions with a wide size range.